One leg of a right triangle is 4m less than the hypotenuse. The other leg is 2m less than the hypotenuse. Find the length of the hypotenuse.
Answer by edjones(8007)

Let c=hypotenuse
a^2+b^2=c^2
(c-2)^2+(c-4)^2=c^2
c^2-4c+4+c^2-8c+16=c^2
c^2-12c+20=0
(c-2)(c-10)=0
c=10 m length of the hypotenuse.
.
Ed